OpenStudy (anonymous):

A cylinder has a height of 14 centimeters and its circle bases have a radius of 10 centimeters. Find the surface area of the cylinder.

OpenStudy (anonymous):

Do you know the formula for the surface area of a cylinder?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

http://www.math.com/tables/geometry/surfareas.htm

OpenStudy (anonymous):

The only two values you need are r and h. "r" is your radius of 10 cm and "h" is your height of 14 cm. Plug and chug.

OpenStudy (anonymous):

dk how

OpenStudy (anonymous):

I gave you the website that lists surface areas for all kinds of shapes. Find where it says "surface area of a cylinder" and the formula's right there.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

If you get stuck just post what you have and we'll help.

OpenStudy (anonymous):

So did you find the formula?

OpenStudy (anonymous):

2 pi r 2 + 2 pi r h

OpenStudy (anonymous):

Ok, and you know your r and h, right? And "pi" is just 3.14, right? Or if you have a scientific calculator it should just have the "pi" button on there. So plug in r=10 for the r in that formula and plug in h=14 for the h and multiply it all out. (2 times pi times r times 2) + (2 times pi times r times h). Do you have trouble multiplying?

OpenStudy (anonymous):

no but im soo confused

OpenStudy (anonymous):

Do you not have any previous experience plugging in numbers into a formula? Ok here it is: \[2*\pi*r ^{2} + 2*\pi*r*h\]\[2*\pi*10^{2}+2*\pi*10*14\]\[(2*3.14*100) +(2*3.14*140)\]

OpenStudy (anonymous):

Do you have a calculator handy? If so, do the first parentheses first: 2 times 3.14 times 100 is equal to 628. The second parentheses is 2 times 3.14 times 140 which is equal to 879.2 so add those two parentheses together: 628 + 879.2 = ?

OpenStudy (anonymous):

i have no cal

OpenStudy (anonymous):

Ok well I did the calculator part for you anyway. The final result is in my last sentence -- you can add those two numbers just by writing them out. What do you get when you add 628 and 879.2? That's the surface area of your cylinder.

OpenStudy (anonymous):

1507.2

OpenStudy (anonymous):

Yep. Good job.

OpenStudy (anonymous):

thx